OpenStudy (anonymous):

Find the exact area of the circle with the given radius. r = 4.5 m A. 4.5π m² B. 9π m² C. 20.25π m² D. 81π m²

OpenStudy (***[isuru]***):

\[A = \pi \ r^{2}\]

OpenStudy (anonymous):

A = (pi)r^2 = (pi)4.5^2 = 20.25(pi)m^2 (C)

OpenStudy (***[isuru]***):

this is the equation to find Area... and here r is the radius.... in ur question u got 4.5 m as the radius...

OpenStudy (anonymous):

yeah ok but i answered his question

OpenStudy (***[isuru]***):

\[A = \ \pi \ (4.5)^{2}\]\[A = \pi \ 20.25 m^{2}\]

OpenStudy (anonymous):

Find the exact area of the circle with the given diameter. d = 2.2 m A. 1.21π m² B. 2.2π m² C. 4.4π m² D. 4.84π m²

OpenStudy (***[isuru]***):

\[radius = \frac{ diameter }{ 2 }\]

OpenStudy (***[isuru]***):

u can find the radius using the above method... then plug the value f the radius to the equation.. \[A = \ \pi \ r^{2}\]